4)Engineers must consider the diameters of heads when designing helmets. The company researchers have determined that the population of potential clientele have head diameters that are normally distributed with a mean of 5.8-in and a standard deviation of 1-in. Due to financial constraints, the helmets will be designed to fit all men except those with head diameters that are in the smallest 0.9% or largest 0.9%.

a)What is the minimum head diameter that will fit the clientele?
min =

b)What is the maximum head diameter that will fit the clientele?
max =

(Enter your answer as a number accurate to 1 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.)
Answer #1

In this

mean = \mu = 5.8 in.

standard deviation = \sigma = 1 in.

a.) the smallest 0.9%

Using standard normal table,

P(Z < z) = 0.9%

= P(Z < z ) = 0.009

= P(Z < -2.36 ) = 0.009

z = -2.362

x = z * \sigma + \mu

x = -2.362 * 1 + 5.8

x = 3.438 in.

Minimum = 3.4 in.

b.) the largest 0.9%

P(Z > z) = 0.9%

= 1 - P(Z < z) = 0.009

= P(Z < z) = 1 - 0.009

= P(Z < z ) = 0.991

= P(Z < 2.362 ) = 0.991  

z = 2.362

x = z * \sigma + \mu

x = 2.362 * 1 + 5.8

x = 8.162 in.

maximum = 8.2 in.